line x/6+y/8=0 intersect the x and y axis at M and N respectively.If the coordinate of point p lying inside the triangle OMN where O is origin.are (a,b,) such that area of the triangle POM,PMN AMD PON are equal find (1) the coordinate of P (2) The radius of the circle iscribed opposite to the angle N

Grade:11

1 Answers

Arun
6 years ago
Dear Kamal
 
M= (6,0)
N = (0, 8)
O= (0, 0)
P will be the centroid
Hence P = (6/3, 8/3) = (2, 8/3)
 
Also
r1\Delta/(s-a)
Where \Delta = 24, s = 24/2 = 12 , a = 6
Hence
r1 = 24/(12-6) = 24/6 = 4
 
Regards
Arun (askIITians forum expert)